OpenStudy (anonymous):

i jas wanna practice some questns on differentiation...................a test would be gud plz suggest a website that could fulfill ma needs..........!!!!!!

OpenStudy (dumbcow):

http://www.math.ucdavis.edu/~kouba/ProblemsList.html this has some good ones for chain rule, product rule, quotient rule...

OpenStudy (anonymous):

\[\text{Differentiate } (x+1)^{(x+1)^{(x+1)}}\]

OpenStudy (anonymous):

\[\text{Differentiate (nicely) } \sqrt{\frac{1}{(x^2+1)(3x-4)^4}}\] \[\text{Differentiate (nicely) } \frac{x}{(x-1)(x+3)^2(x^2-1)}\] \[\text{Differentiate } a^x + b^x\] \[\text{Differentiate } x^{\sin x}\] \[\text{Differentiate } (\sin x)^{x}\] \[\text{Differentiate } (x +x^2)^{x}\] \[\text{Differentiate } (x +1)^{x+1}\] \[\text{Differentiate } x^{x^{x}}\]

OpenStudy (anonymous):

Ugh, annoyingly, that site linked is actually pretty good, so (most) of my problems are rendered moot. Grrrrr.

OpenStudy (dumbcow):

no they are good, very challenging or intimidating i should say

OpenStudy (anonymous):

thnx ...newton

OpenStudy (anonymous):

Indeed, 'intimidating' would be better. If you don't see the 'trick', then they are very hard. However, they are unfortunately all very similar, and one section on that site (won't tell 'cause it'll spoil) has lots like it - and harder ones.

OpenStudy (dumbcow):

yep its a great study tool

OpenStudy (anonymous):

wat is 'intimidating'............ new term 4 me!!!!!!

OpenStudy (anonymous):

Scary/frightening . But slightly different.